# When in Doubt, Deliberate: Confidence-Based Routing to Expert Debate for Sexism Detection

### v1

@@ Abstract
Sexist content online increasingly appears in subtle, context-dependent forms that evade traditional detection methods. Its interpretation often depends on overlapping linguistic, psychological, legal, and cultural dimensions, which produce mixed and sometimes contradictory signals, even in annotated datasets. These inconsistencies, combined with label scarcity and class imbalance, result in unstable decision boundaries and cause fine-tuned models to overlook subtler, underrepresented forms of harm. Together, these limitations point to the need for a design that explicitly addresses the combined effects of (i) underrepresentation, (ii) noise, and (iii) conceptual ambiguity in both data and model predictions. To address these challenges, we propose a two-stage framework that unifies (i) targeted training procedures to adapt supervision to scarce and noisy data with (ii) selective, reasoning-based inference to handle ambiguous or borderline cases. Our training setup applies class-balanced focal loss, class-aware batching, and post-hoc threshold calibration to mitigate label imbalance and noisy supervision. At inference time, a dynamic routing mechanism classifies high-confidence cases directly and escalates uncertain instances to a novel Collaborative Expert Judgment (CEJ) module, which prompts multiple personas and consolidates their reasoning through a judge model.Our approach achieves state-of-the-art results across several benchmarks, with a +2.72% improvement in F1 on the EXIST 2025 Task 1.1, and a gains of +4.48% and +1.30% on the EDOS Tasks A and B, respectively.
When in Doubt, Deliberate: Confidence-Based Routing to Expert Debate for Sexism Detection
@@ 1 Introduction
Figure 1: Overview of the proposed routing mechanism. Predictions with low confidence from the fine-tuned model are escalated to a Collaborative Expert Judgment (CEJ) module, where multiple LLM-based personas engage in structured debate, before the final decision.
@@
To address these challenges, we introduce a two-stage framework that combines (i) targeted training procedures to adapt supervision to scarce and noisy data, with (ii) selective, reasoning-based inference to handle ambiguous cases. This joint design enables the specialised model to make efficient decisions when possible, while engaging with finer contextual reasoning when necessary. During (i) training, we mitigate long-tailed label distributions and inconsistent annotation through three well-known training strategies. These have never been adapted into an integrated framework for such tasks: the class-balanced focal loss (CB-Focal) (ce), which amplifies the contribution of minority and difficult examples, and the class-aware batching, which ensures uniform exposure to underrepresented classes during optimization. We further employ post-hoc threshold calibration to stabilize the decision boundaries of the specialist classifier and improve robustness under noisy supervision. At (ii) inference time, we deploy confidence-aware dynamic routing, which filters the model predictions by difficulty: high-confidence samples are directly classified, while uncertain cases are escalated to a Collaborative Expert Judgment (CEJ) module. The CEJ module implements a structured reasoning pipeline composed of multiple personas, such as a linguist, a gender expert, a psychologist, and an everyday speaker, each prompted to analyze the same input from distinct perspectives. Their deliberations are subsequently synthesized by a judge model, which integrates these perspectives into a final, interpretable decision (Figure 1). The aim is not to faithfully replicate human interaction, but rather to approximate its structure to activate the reasoning and knowledge that instruction-tuned LLMs already encode, yet rarely employ in single-step classification, as evidenced by the limited predictive power of zero-shot baselines edos23; exist25; khan2025explaining.
@@
Our contributions can be summarized as follow:
@@ 2 Related Work
Our work is aligned to at least two research line on sexissm detection and persona-based analyses.
@@
Annotator disagreement, often shaped by cultural perspectives, complicates sexism detection davani2022disagreement. Initiatives like the EXIST 2024 shared task encouraged methods robust to inconsistent annotations. Successful approaches leveraged transformer-based models such as DistilBERT sanh2019distilbert, DeBERTa he2020deberta, RoBERTa liu2019roberta, and hybrid ensembles that utilized data augmentation and multi-task learning. When it comes to LLMs, previous studies have employed adversarial evaluations and data augmentation to enhance detection accuracy. For example, samory2021call highlighted standard model vulnerabilities through psychologically-grounded adversarial datasets, emphasizing the need for psychologically informed robustness. Similarly, khan2025explaining addressed annotator disagreements through definition-driven data augmentation and ensemble methods, improving data-level robustness.
Beyond explicit abuse, a growing challenge lies in detecting nuanced or implicit forms of hate. The work by zeng2025sheeps show that LLMs often fail to recognize metaphorical expressions of hate speech, highlighting substantial gaps in current moderation capabilities. Their findings emphasize that even state-of-the-art models struggle when hateful intent is obscured through indirect language, a phenomenon that is particularly relevant for sexist discourse. Building on this, the authors in fasching2025model demonstrate that hate speech detection outcomes vary significantly across LLM-based systems, underscoring how moderation decisions remain model-dependent and inconsistent. Our work directly responds to these challenges by integrating confidence-aware routing and multi-expert arbitration, ensuring that inconsistent or ambiguous cases are addressed through structured reasoning and synthesis.
@@ 3.3 Collaborative Expert Judgment (CEJ)
# Variant Model ICM ICM-Norm F1_YES F1 Zero-shot Baselines 1 LLaMA-3.3-70B-Instruct 0.3002 0.6509 0.6868 0.7145 2 Cogito-70B 0.3656 0.6837 0.7109 0.7348 3 Qwen2.5-72B-Instruct 0.3765 0.6892 0.7112 0.7384 4 Cogito-70B (reasoning) 0.3772 0.6896 0.7143 0.7384 EXIST 2025 Task 1.1 Leaderboard 5 BERT nowakowski2025exist 0.5727 0.7878 0.7802 6 XLM-RoBERTa pan2025exist 0.5799 0.7915 0.7824 7 Dual-Transformer Fusion Network khan2025exist 0.5806 0.7918 0.7837 8 DeepSeek-R1-Distill-Llama-8B villarreal2025exist 0.6127 0.8079 0.7945 9 Ensemble Approach c:22 0.6249 0.8141 0.7991 10 XLM-RoBERTa c:21 0.6297 0.8165 0.7996 11 LLaMA-3.1-8B-Instruct tian2025marioexist2025simple 0.6774 0.8405 0.8167 Ours 12 $\mathcal{C}_{1}$ LLaMA-3.2-3B 0.5709 0.7913 0.7596 0.7934 13 $\mathcal{C}_{2}$ $\mathcal{C}_{1}$ + Task-specific optimizations 0.6595 0.8315 0.8089 0.8257 14 $\mathcal{C}_{3}$ $\mathcal{C}_{2}$ + Routing to CEJQ→C 0.6831 0.8433 0.8191 0.8348 15 $\mathcal{C}_{4}$ $\mathcal{C}_{2}$ + Routing to CEJL→C 0.6963 0.8500 0.8233 0.8389
Table 1: Performance on the EXIST 2025 Task 1.1 test set. First, zero-shot baselines with generalist LLMs; then, leaderboard submissions at EXIST 2025 Task 1.1; finally, our routing pipeline. “Q” and “L” denote the models used for personas debate (Qwen2.5-72B-Instruct and LLaMA-3.3-70B, respectively), and “C” refers to the judge model Cogito-70B (reasoning mode). Bold indicates the best score, and underline denotes the runner-up.
@@ 3.3.1 Structured Debate and Judge
Summarization. A dedicated summarization agent condense the full debate into a concise synthesis highlighting the main arguments, points of consensus, and unresolved disagreements.
@@
# Variant Model Task A Task B Task C Zero-shot Baselines 1 Qwen-2.5-72B-Instruct 0.6190 0.4246 0.2576 2 LLaMA-3.3-70B-Instruct 0.6421 0.4297 0.2775 3 Cogito-70B 0.6106 0.4214 0.2683 4 Cogito-70B (reasoning) 0.6332 0.4325 0.2826 SemEval 2023 Task 10.A Leaderboard 5 DeBERTa-v3-large + twHIN-BERT-large zhou-2023-pinganlifeinsurance 0.8746 6 RoBERTa-Large + ELECTRA 0.8740 0.7203 0.5487 7 DeBERTa Ensemble 0.8740 Data Augmentation and Ensemble Methods 8 SEFM zhong-etal-2023-uirisc 0.8538 0.6619 0.4641 9 QCon feely-etal-2023-qcon 0.8400 0.6400 0.4700 10 HULAT segura-bedmar-2023-hulat-semeval 0.8298 0.5877 0.4458 11 CSE khan2025explaining 0.8819 0.7243 0.5639 12 DDA khan2025explaining 0.8769 0.7277 0.6018 13 PaLM Ensemble 0.7326 Ours 14 $\mathcal{C}_{1}$ LLaMA-3.2-3B 0.7824 0.5941 0.3717 15 $\mathcal{C}_{2}$ $\mathcal{C}_{1}$ + Task-specific optimizations 0.8986 0.6279 0.4189 16 $\mathcal{C}_{3}$ $\mathcal{C}_{2}$ + Routing to CEJQ→C 0.9195 0.7324 0.5842 17 $\mathcal{C}_{4}$ $\mathcal{C}_{2}$ + Routing to CEJL→C 0.9214 0.7421 0.5904
@@ 4.1 Experimental Setup
Baselines. We evaluate our proposed framework against both fine-tuned and zero-shot LLM baselines. These focuses on state-of-the-art approaches for sexism detection; while an exhaustive exploration of all available LLMs is beyond the scope of this work. As specialized models, we employe LLaMA-3.2-3B, adapted via LoRA fine-tuning (hu2021lora) on both the EDOS and EXIST training datasets. Full task-specific training information are available in Appendix A.2. To evaluate instruction-following and reasoning abilities in a zero-shot setting, we test Qwen2.5-72B-Instruct qwen2025qwen25technicalreport, LLaMA-3.3-70B-Instruct grattafiori2024llama, and Cogito-70B deepcogito_cogito_v2 using simple classification prompts (Appendix A.3). Pre-trained checkpoints were obtained from HuggingFace huggingface. All LLM experiments were executed locally through LangChain langchain and Ollama ollama on a system equipped with three Nvidia A100 GPUs.
@@ 4.2 Results
EXIST Dataset. Results for the the EXIST 2025 Task 1.1 test set, binary classification, are reported in Table 1. The first group of models comprises LLMs evaluated in a zero-shot setting. Among these, Cogito-70B (reasoning mode) achieves the highest F1_YES (0.7143), with closely matched performance from Qwen2.5-72B-Instruct (0.7112) and standard Cogito-70B (0.7109). Overall, all zero-shot models with basic prompts underperform compared to leaderboard approaches, reflecting the limitations of out-of-the-box LLMs for specialized task. The second group reports official leaderboard systems submitted to EXIST 2025 Task 1.1. Here, LLaMA-3.1-8B-Instruct tian2025marioexist2025simple leads in both ICM (0.6774) and F1_YES (0.8167). Transformer fusion architectures and BERT-based systems round out the group, showing the persistence of competitive results of earlier architectures.
The final section of Table 1 details the results of our routing framework. We build upon the LLaMA-3.2-3B speciliased model which achieves performance on par with lower-ranking leaderboard models (F1_YES: 0.7596). Applying targeted optimizations in $\mathcal{C}_{2}$ (multi-dataset training, class-balanced loss, threshold calibration) yields substantial improvements: 0.8089 F1_YES and 0.8257 macro-F1, with results that already surpass many leaderboard submissions. Further gains are realized through confidence-aware dynamic routing. Configuration $\mathcal{C}_{3}$ integrates Qwen2.5-72B-Instruct for persona-based debate, reaching 0.8191 F1_YES, while $\mathcal{C}_{4}$, leveraging LLaMA-3.3-70B for personas and Cogito-70B as judge, achieves the highest scores across all reported metrics. This configuration outperforms the best leaderboard system by 0.66 points on F1_YES and establishes a new benchmark for the task.
@@
The first group report the results of instruction-tuned LLMs evaluated in a zero-shot setting. Performance remains limited across all tasks when using basic prompts. The best macro-F1 for binary classification (Task A) is achieved by LLaMA-3.3-70B-Instruct (0.6421), closely followed by Cogito-70B (reasoning mode, 0.6332) and Qwen2.5-72B-Instruct (0.6190). For fine-grained classification (Task C), all zero-shot models fall below 0.29. Among official leaderboard systems from SemEval 2023 Task 10.A, ensemble methods based on DeBERTa and RoBERTa architectures set a strong baseline, reflecting the effectiveness of large pre-trained transformer models with supervised fine-tuning.
The next group encompasses systems employing data augmentation, model ensembling, or both. Here, CSE khan2025explaining and DDA khan2025explaining achieve notable gains: CSE reaches 0.8819 on Task A and 0.5693 on Task C, while DDA records the highest reported macro-F1 on Task C (0.6018) and competitive results on other tasks.
@@
Table 3 presents class-wise F1 scores for Tasks B and C. CEJ shows an inverse relationship between class frequency and effectiveness: rarer categories benefit most, with gains of +30.7 (Prejudiced Discussions, nn=94) and +17.6 (Threats/Harm, nn=89) in Task B, while the most frequent category (Derogation, nn=454) improves only marginally (+1.9). This pattern holds in Task C, where the smallest classes 3.3 (nn=18) and 3.4 (nn=14) achieve gains of +74.2 and +83.3 points. This suggests that multi-perspective debate compensates for representation bias in underrepresented categories.
@@
Table 4: Prompt ablation (ICM scores, LLaMA-3.3-70B). $\mathcal{P}_{1}$: baseline few-shot prompt with role identity. $\mathcal{P}_{2}$: expands the role into a specialised expert persona. $\mathcal{P}_{3}$: introduces formal sexism definitions aligned with annotation guidelines. $\mathcal{P}_{4}$: adds nuanced and multilingual examples. $\mathcal{P}_{5}$: incorporates refined guidelines informed by error analysis.

### v2

# When in Doubt, Consult: Expert Debate for Sexism Detection via Confidence-Based Routing

@@ Abstract
Sexist content online increasingly appears in subtle, context-dependent forms that evade traditional detection methods. Its interpretation often depends on overlapping linguistic, psychological, legal, and cultural dimensions, which produce mixed and sometimes contradictory signals, even in annotated datasets. These inconsistencies, combined with label scarcity and class imbalance, result in unstable decision boundaries and cause fine-tuned models to overlook subtler, underrepresented forms of harm. Together, these limitations point to the need for a design that explicitly addresses the combined effects of (i) underrepresentation, (ii) noise, and (iii) conceptual ambiguity in both data and model predictions. To address these challenges, we propose a two-stage framework that unifies (i) targeted training procedures to adapt supervision to scarce and noisy data with (ii) selective, reasoning-based inference to handle ambiguous or borderline cases. Our training setup applies class-balanced focal loss, class-aware batching, and post-hoc threshold calibration to mitigate label imbalance and noisy supervision. At inference time, a dynamic routing mechanism classifies high-confidence cases directly and escalates uncertain instances to a novel Collaborative Expert Judgment (CEJ) module, which prompts multiple personas and consolidates their reasoning through a judge model. Our approach achieves state-of-the-art results across several benchmarks, with F1 gains of +4.48% and +1.30% on EDOS Tasks A and B, respectively, and a +2.79% improvement in ICM on EXIST 2025 Task 1.1.
When in Doubt, Consult: Expert Debate for Sexism Detection
via Confidence-Based Routing
@@ 1 Introduction
Figure 1: Overview of the proposed framework.
@@
To address these challenges, we introduce a two-stage framework that combines (i) targeted training procedures to adapt supervision to scarce and noisy data, with (ii) selective, reasoning-based inference to handle ambiguous cases. This joint design enables the specialized model to make efficient decisions when possible, while engaging with finer contextual reasoning when necessary. During (i) training, we mitigate long-tailed label distributions and inconsistent annotation through three well-known training strategies. These have never been adapted into an integrated framework for such tasks: the class-balanced focal loss (CB-Focal) (ce), which amplifies the contribution of minority and difficult examples, and the class-aware batching, which ensures uniform exposure to underrepresented classes during optimization. We further employ post-hoc threshold calibration to stabilize the decision boundaries of the specialist classifier and improve robustness under noisy supervision. At (ii) inference time, we deploy confidence-aware dynamic routing, which filters the model predictions by difficulty: high-confidence samples are directly classified, while uncertain cases are escalated to a Collaborative Expert Judgment (CEJ) module. The CEJ module implements a structured reasoning pipeline composed of multiple personas, such as a linguist, a gender expert, a psychologist, and an everyday speaker, each prompted to analyze the same input from distinct perspectives. Their deliberations are subsequently synthesized by a judge model, which integrates these perspectives into a final, interpretable decision (Figure 1). The aim is not to faithfully replicate human interaction, but rather to approximate its structure to activate the reasoning and knowledge that instruction-tuned LLMs already encode, yet rarely employ in single-step classification, as evidenced by the limited predictive power of zero-shot baselines edos23; exist25; khan2025explaining.
@@
Our contributions can be summarized as follows:
@@ 2 Related Work
Our work is aligned with at least two research line on sexism detection and persona-based analyses.
@@
Annotator disagreement, often shaped by cultural perspectives, complicates sexism detection davani2022disagreement. Initiatives like the EXIST 2024 shared task encouraged methods robust to inconsistent annotations. Successful approaches leveraged transformer-based models such as DistilBERT sanh2019distilbert, DeBERTa he2020deberta, RoBERTa liu2019roberta, and hybrid ensembles that utilized data augmentation and multi-task learning. When it comes to LLMs, previous studies have employed adversarial evaluations and data augmentation to enhance detection accuracy. For example, samory2021call highlighted standard model vulnerabilities through psychologically grounded adversarial datasets, emphasizing the need for psychologically informed robustness. Similarly, khan2025explaining addressed annotator disagreements through definition-driven data augmentation and ensemble methods, improving data-level robustness.
Beyond explicit abuse, a growing challenge lies in detecting nuanced or implicit forms of hate. The work by zeng2025sheeps shows that LLMs often fail to recognize metaphorical expressions of hate speech, highlighting substantial gaps in current moderation capabilities. Their findings emphasize that even state-of-the-art models struggle when hateful intent is obscured through indirect language, a phenomenon that is particularly relevant for sexist discourse. Building on this, the authors in fasching2025model demonstrate that hate speech detection outcomes vary significantly across LLM-based systems, underscoring how moderation decisions remain model-dependent and inconsistent. Our work directly responds to these challenges by integrating confidence-aware routing and multi-expert arbitration, ensuring that inconsistent or ambiguous cases are addressed through structured reasoning and synthesis.
@@ 3.3 Collaborative Expert Judgment (CEJ)
# Variant Model ICM ICM-Norm F1_YES F1 Zero-shot Baselines 1 GPT-5.2 0.3408 0.6705 0.7043 0.7349 2 Qwen2.5-72B-Instruct 0.3765 0.6892 0.7112 0.7384 3 LLaMA-3.3-70B-Instruct 0.3002 0.6509 0.6868 0.7145 4 Cogito-70B 0.3656 0.6837 0.7109 0.7348 5 Cogito-70B (reasoning) 0.3772 0.6896 0.7143 0.7384 EXIST 2025 Task 1.1 Leaderboard 6 BERT nowakowski2025exist 0.5727 0.7878 0.7802 7 XLM-RoBERTa pan2025exist 0.5799 0.7915 0.7824 8 Dual-Transformer Fusion Network khan2025exist 0.5806 0.7918 0.7837 9 DeepSeek-R1-Distill-Llama-8B villarreal2025exist 0.6127 0.8079 0.7945 10 Ensemble Approach c:22 0.6249 0.8141 0.7991 11 XLM-RoBERTa c:21 0.6297 0.8165 0.7996 12 LLaMA-3.1-8B-Instruct tian2025marioexist2025simple 0.6774 0.8405 0.8167 Ours 13 $\mathcal{C}_{1}$ LLaMA-3.2-3B 0.5709 0.7913 0.7596 0.7934 14 $\mathcal{C}_{2}$ $\mathcal{C}_{1}$ + Task-specific optimizations 0.6595 0.8315 0.8089 0.8257 15 $\mathcal{C}_{3}$ $\mathcal{C}_{2}$ + Routing to CEJQ→C 0.6831 0.8433 0.8191 0.8348 16 $\mathcal{C}_{4}$ $\mathcal{C}_{2}$ + Routing to CEJL→C 0.6963 0.8500 0.8233 0.8389
Table 1: Results on the EXIST 2025 Task 1.1 test set with zero-shot baselines, leaderboard submissions, and our approach. "Q" and "L" denote the models used for personas debate (Qwen2.5-72B-Instruct and LLaMA-3.3-70B, respectively), and "C" refers to the judge model Cogito-70B (reasoning mode). Bold = best score, underline = runner-up.
@@ 3.3.1 Structured Debate and Judge
Summarization. A dedicated summarization agent condenses the full debate into a concise synthesis highlighting the main arguments, points of consensus, and unresolved disagreements.
@@
# Variant Model Task A Task B Task C Zero-shot Baselines 1 GPT-5.2 0.7306 0.4543 0.3381 2 Qwen-2.5-72B-Instruct 0.6190 0.4246 0.2576 3 LLaMA-3.3-70B-Instruct 0.6421 0.4297 0.2775 4 Cogito-70B 0.6106 0.4214 0.2683 5 Cogito-70B (reasoning) 0.6332 0.4325 0.2826 SemEval 2023 Task 10.A Leaderboard 6 DeBERTa-v3-large + twHIN-BERT-large zhou-2023-pinganlifeinsurance 0.8746 7 RoBERTa-Large + ELECTRA 0.8740 0.7203 0.5487 8 DeBERTa Ensemble 0.8740 Data Augmentation and Ensemble Methods 9 SEFM zhong-etal-2023-uirisc 0.8538 0.6619 0.4641 10 QCon feely-etal-2023-qcon 0.8400 0.6400 0.4700 11 HULAT segura-bedmar-2023-hulat-semeval 0.8298 0.5877 0.4458 12 CSE khan2025explaining 0.8819 0.7243 0.5639 13 DDA khan2025explaining 0.8769 0.7277 0.6018 14 PaLM Ensemble 0.7326 Ours 15 $\mathcal{C}_{1}$ LLaMA-3.2-3B 0.7824 0.5941 0.3717 16 $\mathcal{C}_{2}$ $\mathcal{C}_{1}$ + Task-specific optimizations 0.8986 0.6279 0.4189 17 $\mathcal{C}_{3}$ $\mathcal{C}_{2}$ + Routing to CEJQ→C 0.9195 0.7324 0.5842 18 $\mathcal{C}_{4}$ $\mathcal{C}_{2}$ + Routing to CEJL→C 0.9214 0.7421 0.5904
@@ 4.1 Experimental Setup
Baselines. We evaluate our proposed framework against both fine-tuned and zero-shot LLM baselines. These focus on state-of-the-art approaches for sexism detection; while an exhaustive exploration of all available LLMs is beyond the scope of this work. As specialized models, we employ LLaMA-3.2-3B, adapted via LoRA fine-tuning (hu2021lora) on both the EDOS and EXIST training datasets. Full task-specific training information is available in Appendix A.2. To evaluate instruction-following and reasoning abilities in a zero-shot setting, we test Qwen2.5-72B-Instruct qwen2025qwen25technicalreport, LLaMA-3.3-70B-Instruct grattafiori2024llama, GPT-5.2 openai2025gpt5, and Cogito-70B deepcogito_cogito_v2 using simple classification prompts (Appendix A.3). Pre-trained checkpoints were obtained from HuggingFace huggingface. All LLM experiments were executed locally through LangChain langchain and Ollama ollama on a system equipped with three Nvidia A100 GPUs.
@@ 4.2 Results
EXIST Dataset. Results for the EXIST 2025 Task 1.1 test set, binary classification, are reported in Table 1. The first group of models comprises LLMs evaluated in a zero-shot setting. Among these, Cogito-70B (reasoning mode) achieves the highest F1_YES (0.7143), outperforming Qwen2.5-72B-Instruct (0.7112), standard Cogito-70B (0.7109), and state-of-the-art GPT-5.2 (0.7043). Overall, all zero-shot models with basic prompts underperform compared to leaderboard approaches, reflecting the limitations of out-of-the-box LLMs for specialized tasks.
The second group reports official leaderboard systems submitted to EXIST 2025 Task 1.1. Here, LLaMA-3.1-8B-Instruct tian2025marioexist2025simple leads in both ICM (0.6774) and F1_YES (0.8167). Transformer fusion architectures and BERT-based systems round out the group, showing the persistence of competitive results of earlier architectures.
The final section of Table 1 details the results of our routing framework. We build upon the LLaMA-3.2-3B speciliazed model which achieves performance on par with lower-ranking leaderboard models (F1_YES: 0.7596). Applying targeted optimizations in $\mathcal{C}_{2}$ (multi-dataset training, class-balanced loss, threshold calibration) yields substantial improvements: 0.8089 F1_YES and 0.8257 macro-F1, with results that already surpass many leaderboard submissions. Further gains are realized through confidence-aware dynamic routing. Configuration $\mathcal{C}_{3}$ integrates Qwen2.5-72B-Instruct for persona-based debate, reaching 0.8191 F1_YES, while $\mathcal{C}_{4}$, leveraging LLaMA-3.3-70B for personas and Cogito-70B as judge, achieves the highest scores across all reported metrics. This configuration outperforms the best leaderboard system by 0.66 points on F1_YES and establishes a new benchmark for the task.
@@
The first group reports the results of LLMs evaluated in a zero-shot setting. Performance remains limited across all tasks when using basic prompts. GPT-5.2 achieves the highest performance across all tasks (Task A: 0.7306, Task B: 0.4543, Task C: 0.3381), followed by LLaMA-3.3-70B-Instruct (0.6421). Cogito-70B in reasoning mode consistently outperforms its standard variant (+2.26 pp in Task A). However, all zero-shot models struggle with fine-grained classification (Task C < 0.34).
Among official leaderboard systems from SemEval 2023 Task 10.A, ensemble methods based on DeBERTa and RoBERTa architectures set a strong baseline, reflecting the effectiveness of large pre-trained transformer models with supervised fine-tuning.
The next group encompasses systems employing data augmentation, model ensembling, or both. Here, CSE khan2025explaining and DDA khan2025explaining achieve notable gains: CSE reaches 0.8819 on Task A and 0.5639 on Task C, while DDA records the highest reported macro-F1 on Task C (0.6018) and competitive results on other tasks.
@@
Table 3 presents class-wise F1 scores for Tasks B and C. CEJ shows an inverse relationship between class frequency and effectiveness: rarer categories benefit most, with gains of +30.7 (Prejudiced Discussions, nn=94) and +17.6 (Threats/Harm, nn=89) in Task B. This pattern holds in Task C, where the smallest classes 3.3 (nn=18) and 3.4 (nn=14) achieve gains of +74.2 and +83.3 points. This suggests that multi-perspective debate compensates for representation bias in underrepresented categories.
@@
Table 4: Prompt ablation (ICM scores, LLaMA-3.3-70B). $\mathcal{P}_{1}$: baseline few-shot with role identity. $\mathcal{P}_{2}$: specialized expert persona. $\mathcal{P}_{3}$: formal sexism definitions. $\mathcal{P}_{4}$: nuanced multilingual examples. $\mathcal{P}_{5}$: refined guidelines from error analysis.